# Towards Next-Generation Urban Decision Support Systems through AI-Powered Generation of Scientific Ontology using Large Language Models - A Case in Optimizing Intermodal Freight Transportation

### v1

@@ 4.2 Ontology-driven Decision Optimization
Figure 6: Decision support workflow that is driven by the scenario-based ontology created using the LLM-powered autonomous workflow.
The case study presented in this section utilizes a subset of FAF, FTOT, and AFS BTS data to establish a use case for identifying the optimal combination of shipping modes and routes from Nashville to New Orleans. Traditional manual approaches to decision optimization often demand extensive domain expertise regarding the dynamics of inter-modal freight transportation events, as well as data- and simulation-driven insights for quantifying various decision support criteria, such as operational costs and emissions.
@@
Employing our prototyping methodology for creating scenario-based ontology, we have successfully developed a digital representation of knowledge sourced from the technical manuals of major intermodal freight transportation datasets and simulations. This digital knowledge representation is exported into an OWL format, furnishing an ontology-driven pipeline to delineate the decision support workflow and to structure data and simulation outputs within the domain context. Our ontology consists of three major components, as the following:
@@
    Graph Representation is a graph data model that captures the topological properties (connectivity) of the freight transportation network. It is derived by the LLM using information from the FAF dataset manual. The model represents various modes and routes of freight transportation between key origin-destination pairs, which are represented by US cities that are major logistic and transportation hubs. This graph is primarily used to navigate between any user-defined origin-destination pairs.
@@
    Data Sources comprise all the necessary multi-domain datasets required to provide data- and simulation-driven insights, complete with essential decision variables. These insights aid in making informed decisions aimed at optimizing operational costs and delivery times, while also reducing fuel consumption and GHG emission. These data sources are organized through a scenario-based ontology and are mapped to corresponding network entities (nodes and edges). Subsequently, they are ingested into a relational database, utilizing a data schema derived from the ontology to organize the data to enable rapid data retrieval for decision support purposes.
@@
    The “Network Traverse” refers to the systematic process of navigating through a transportation network to identify all possible combinations of routes and modes between an origin-destination pair. This process involves assessing various transportation modes—such as road, rail, and river—and their interconnections, along with transit locations represented as intermediate nodes within the FTOT network. Each mode and route combination is characterized by a series of nodes that signify the origin, destination, and intermodal transits, connected by edges that represent the transitions between modes. This comprehensive approach ensures a thorough exploration of the network, facilitating optimal route and mode selection.
@@
Our methodology capitalizes on the capabilities of LLMs to automate the digitization of domain knowledge encapsulated in text and figures within technical manuals of urban datasets, simulations, and research articles. This knowledge is converted into interoperable ontology formats, which underpin the development of data models. These models structure essential data and simulation outputs to support decision-making processes. The lookup table, generated in real-time from the ontology-driven data model, plays a pivotal role. Together, they form crucial components for the development of urban decision support systems or digital twin applications, enhancing the efficiency and effectiveness of urban planning and management.
@@ 5 Limitation and Future Work
    Without precisely engineered prompts, the ChatGPT API may exhibit sensitivity, leading to subtle variability and inconsistent results across different iterations. The robustness of the model can be enhanced by employing more advanced techniques, such as fine-tuning and prompt tuning. These methods adapt pre-trained LLMs to generate ontologies using specific datasets or under particular conditions, thereby minimizing potential variability in the results from different model runs.
@@
We aim to address these limitations through our future work.
@@ 6 Conclusion
In this study, we introduced an AI-powered methodology designed to automate the generation of scenario-based scientific ontologies by harnessing the text understanding and reasoning capabilities of LLMs. This methodology was implemented into a prototype autonomous agent using Python NLP libraries and the ChatGPT API. Our primary goal is to support the development of urban decision support systems by automating time-consuming and labor-intensive tasks, which typically require substantial domain knowledge. Our approach facilitates the integration of data resources and simulations to enhance informed decision-making.
We tested our methodology by generating the well-known Pizza Ontology and compared the AI-generated outputs with the expert-created Pizza Ontology, validating the reliability and feasibility of our approach. Through a case study, we demonstrated how our methodology could be applied to real-world urban data, such as the FAF and FTOT datasets, to aid in the development of an ontology-driven urban decision support system. Although still in its early stages, our approach showcases the potential and feasibility of using an autonomous agent to generate ontologies that support scientific software and database development. This initiative sheds light on the next generation of decision support systems that leverage large-scale urban data to optimize complex urban systems.
With significant promise for achieving autonomous decision support systems, we believe our methodology can be further extended and applied to the creation of various domain ontologies, thereby lowering technical barriers and enhancing the capability to solve complex urban system challenges. Additionally, our methods can be leveraged to automate bibliometric analyses, focusing on the extraction of key domain concepts. Future work of this study includes applying more recent model tuning techniques, such as the prompt tuning and fine tuning, to further instruct LLM on generating more specialized ontology, as well as software modules and codes that serve as the critical components of a urban decision support system.

### v3

@@ 5.5 Decision Support Strategies - Formulation
Figure 7: Conceptual decision support workflow that is driven by the scenario-based ontology created using the LLM-powered autonomous workflow.
@@
By employing our prototyping methodology for creating scenario-based ontology, we have successfully developed a digital representation of knowledge sourced from the technical manuals of major intermodal freight transportation datasets and simulations. This digital knowledge representation is exported into an OWL format, providing an ontology-driven pipeline to outline the decision support workflow and structure data and simulation outputs within the domain context. Our ontology consists of three major components, as the following:
@@
    Graph Representation is a graph data model that captures the topological properties (connectivity) of the freight transportation network. It is derived by the LLM using information from the FAF dataset manual. The model represents various modes and routes of freight transportation between key OD pairs, which are represented by US cities that are major logistic and transportation hubs. This graph is primarily used to navigate between any user-defined origin-destination pairs.
@@
    Data Sources comprise all the necessary multi-domain datasets required to provide data- and simulation-driven insights, complete with essential decision variables. These insights aid in making informed decisions aimed at optimizing operational costs and delivery times while also reducing fuel consumption and GHG emissions. These data sources are organized through a scenario-based ontology and are mapped to corresponding network entities (nodes and edges). Subsequently, they are ingested into a relational database, utilizing a data schema derived from the ontology to organize the data to enable rapid data retrieval for decision support purposes.
@@
    The “Network Traverse” refers to the systematic process of navigating through a transportation network to identify all possible combinations of routes and modes between an OD pair. This process involves assessing various transportation modes—such as road, rail, and river—and their interconnections, along with transit locations represented as intermediate nodes within the FTOT network. Each mode and route combination is characterized by a series of nodes that signify the origin, destination, and intermodal transits, connected by edges that represent the transitions between modes. This comprehensive approach ensures a thorough exploration of the network, facilitating optimal route and mode selection.
@@
Our methodology leverages the capabilities of LLMs to automate the digitization of domain knowledge encapsulated in text and figures within technical manuals, urban datasets, simulations, and research articles. This knowledge is transformed into interoperable ontology formats, which serve as the foundation for developing data models. These models structure essential data and simulation outputs to support decision-making processes. The real-time generated lookup table from the ontology-driven data model is a critical component. Together, these elements form essential parts of urban decision support systems or digital twin applications, enhancing the efficiency and effectiveness of urban planning and management.
5.6 Decision Support Strategies - Implementation
We implemented our proposed methodology using the AI-generated ontology and its knowledge graph to provide data source descriptions, guiding the design and queries of databases that are essential components of the data provisioning system. Figure 8 demonstrates the technical pipeline, outlining the processes involved in supporting decision-making to reduce GHG emissions, operational costs, fuel consumption, and shipment time using structured data and simulation outputs. The first step in the technical pipeline is defining OD pairs and optimization requirements by the target users. These inputs are directly fed into the neo4j graph database and the AI-generated ontology and knowledge graph to configure the optimization scenario. Based on the OD, the graph database conducts a shortest path query that identifies a list of possible combinations of modes and routes between the OD based on transport distance. Example queries of the graph database operations are provided in Figure 8. Through the query, each combination of mode and route will be identified, specifying the mode with fuel types, node IDs representing hubs and intermodal transit points, and edge IDs describing the attributes of the transportation path (e.g., rail, road, and river).
Figure 8: Technical implementation of the decision support workflow utilizes the proposed ontology-driven methods to guide the design and queries of PostgreSQL and neo4j databases.
By configuring the ontology, we define the constraints and requirements of user scenarios, guiding the inclusion and exclusion of specific decision criteria. In the demonstration, we showcase a scenario aimed at identifying the most optimized shipping solution between Nashville and New Orleans, as conceptually depicted in Figure 7. This scenario will be optimized based on criteria including GHG emissions, shipment distance, and shipment time. Based on the query results from the graph database, the AI-generated ontology defines the decision support metrics and provides data and simulation descriptions through its knowledge graph. This guides the discovery of essential decision metrics stored in the PostgreSQL relational database. This setup allows for the future inclusion of more ontology-driven parameters, such as route disruptions, free-speed travel, and capacity. For this test scenario, we focus solely on distance, node vicinity, metrics, and constraints.
Once the variables characterizing decision support metrics are retrieved for each combination of modes and routes, a Python script aggregates these metrics to calculate the total amount of GHG emissions, time, and cost. The various combinations identified through the graph database, along with their defined metrics, are then used as inputs for the GUROBI optimization engine.
@@ 6 Limitation and Future Work
    Without precisely engineered prompts, the ChatGPT API may exhibit sensitivity, leading to subtle variability and inconsistent results across different iterations. The robustness of the model can be enhanced by employing more advanced techniques, such as fine-tuning. These methods adapt pre-trained LLMs to generate ontologies using specific datasets or under particular conditions, thereby minimizing potential variability in the results from different model runs.
@@
  3. 3.
    Potential hallucinations of LLMs could jeopardize the accuracy and reliability of AI-generated ontologies and knowledge graphs. To minimize these hallucinations, previous studies have proposed matching techniques that utilize vocabularies defined in existing domain ontologies to validate AI-generated content (Caufield et al., 2024). Additionally, leveraging external knowledge bases through the recent emergence of RAG can also help reduce hallucinations. Our method demonstrates a qualitative analysis that logically addresses the competency questions proposed through SPARQL queries.
We aim to address these limitations through our future work by deploying and fine-tuning a local pre-trained model using Ollama.
@@ 7 Conclusion
In this study, we introduced an AI-powered methodology designed to automate the generation of scenario-based scientific ontologies by leveraging the text understanding and reasoning capabilities of LLMs. This methodology was implemented in a prototype autonomous agent using Python NLP libraries and the ChatGPT API. Our primary goal is to support the development of urban decision support systems by automating time-consuming and labor-intensive tasks that usually require extensive domain knowledge. Our automated approach facilitates the integration of data resources and simulations, thereby enhancing informed decision-making.
We tested our methodology by generating a partial segment of the Pizza Ontology and compared the AI-generated outputs with the expert-created Pizza Ontology, validating the reliability and feasibility of our approach. Through a case study, we demonstrated how our methodology could be applied to real-world urban data, such as the FAF and FTOT datasets, to aid in the development of an ontology-driven urban decision support system. Although still in its early stages, our approach showcases the potential and feasibility of using an autonomous workflow to generate ontologies that support scientific software and database development, as well as guide the formation of decision support strategies. This initiative highlights the next generation of decision support systems and digital twins that leverage large-scale urban data to optimize complex urban systems.
With significant promise for achieving autonomous decision support systems, we believe our methodology can be further extended and applied to creating various scenario-based ontologies. This extension could lower technical barriers and enhance the ability to address complex urban system challenges. Future work will focus on applying fine-tuning techniques to instruct LLMs on generating more specialized ontologies, as well as developing software modules and codes that serve as critical components of an urban decision support system.